OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x - 9. 7 -7 11 -11 @sodukogirl13

OpenStudy (anonymous):

-11

OpenStudy (anonymous):

5x+2=4x-9 -2 -2 5x=4x-11 -4x -4x x=-11